The length of a rectangle is
5
inches more than
3
times the width. The perimeter is
66
inches. Find the length and the width

Answer by ikleyn(52781)

.

   The width + the length = 66/2 = 33


      w      + (3w + 5)          = 33


      w      +  3w               = 33-5

          4w                     = 28

           w                     = 28/4 = 7.


ANSWER.  The width is 7 inches.  The length is  3*7+5 = 26 inches.

Solved.

--------------

Hey, by respecting the tutors, can you change the formatting of your posts in the future, P L E A S E ?